OpenStudy (anonymous):

The height of a ball tossed vertically up from 15 m with speed of 10m/s described by h(t) =15+10t-5t time in sec after ball thrown and h height above ground How many sec before ball falls to ground What is max height Pls help cheers

OpenStudy (jack1):

h(t) =15 + 10t - 5t should one of these be t^2 ...?

OpenStudy (anonymous):

if its-5t^2+10t+15, then the ball will be in the air for 3 seconds and reach a max height of 20

OpenStudy (anonymous):

yes one is 5t squared

OpenStudy (anonymous):

Why did you put the 15 outside the brackets and where did the 4 come from? Cheers

OpenStudy (anonymous):

at t=1 ud be at -5+10+15 = 20, vertex is at (1,20)

OpenStudy (the_fizicx99):

-5t^2+10t+15 Vertex: \(\ \sf x = \dfrac{-b}{2a} \) a = -5 b = 10 \(\ \sf x = \dfrac{-10}{-10} \implies 1\) Now substitute that into the function: h(t) = -5t^2+10t+15 and that'll be the "y" part of the coordinate. you already have the "x," hence the "x = " part. Therefore (1, ?) is the vertex ~ Finish it